# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ 5 Application to classification
Here we consider the classification case $\mathcal{Y}=[K]$, apply our results and provide some new ones. We deal here both with the iid model and the class-conditional model.
5.1 Post-processing
In classification, any selective prediction set procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is post-processed by setting, for $i\in\mathcal{S}$, $\mathcal{C}_{n+i}=\arg\min_{k\in[K]}\{S_{k}(X_{n+i})\}$ whenever $\mathcal{C}_{n+i}=\emptyset$ (that is, if empty take the smallest non-conformity score). Clearly, this operation can only decrease the FCP while it can only increase the adjusted power, so it should always be preferred in the classification case. In this paper, InfoSP and InfoSCOP always refer to the post-processed procedures in the classification case.
@@ 5.2 Choosing the appropriate $p$-value collection in classification
Thus, for the iid model, we can in principle use either full-calibrated $p$-values or class-calibrated $p$-values. In the applications we consider next, it appears that using full-calibrated $p$-values in InfoSP is best. We support this claim by theory for non-trivial classification in § 5.4, and by numerical experiments in § 5.3, § 5.4, and § F. We note that for the initial selection step in InfoSCOP, class-calibrated $p$-values can be useful, as demonstrated in § 5.3.
For the class-conditional model, there can be a label shift from the calibration to the test sample. So the full-calibrated $p$-values are not valid, and class-calibrated $p$-values must be used. In § B we consider more generally weighted class-calibrated $p$-values, where the weights are functions of estimators of the proportion of labels in each class in the test sample.
@@ 5.3 An illustrative example: prediction sets excluding a null class
Suppose the analyst is interested in reporting prediction sets that exclude a null class, say class $y_{0}=1$ (see first item of Examples 1.2 and 2.2). We consider the following novel procedures, in addition to the naive procedure using the classic conformal procedure, that reports $\mathcal{C}^{\alpha}_{n+i}$ only if $\mathcal{C}^{\alpha}_{n+i}$ does not contain the null class:
  • InfoSP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$.
  • InfoSCOP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})$, with initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ being the BH procedure applied to the class-calibrated adaptive $p$-value family $(\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}},i\in[r+1,n+m])$, given by
    $\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}}=\hat{\pi}_{1}\tilde{p}^{(1)}_{i},\>i\in[r+1,n+m],$
    using $\{(X_{j},Y_{j}),j\in[r]\}$ and $\{(X_{j},Y_{j}),j\in[r+1,n+m]\}$ as calibration and test samples, respectively, and with calibrated-based estimator $\hat{\pi}_{1}=\frac{\sum_{i=1}^{r}{\mathbf{1}\{Y_{i}=1\}}+1}{r+1}$.
  • InfoSP on class-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}})$.
Note that for InfoSCOP above uses class-calibrated $p$-values for initial selection (because these are better to detect examples from the non-null class), and full-calibrated $p$-values on the selected examples from $[r+1,n+m]$ in the second step (because these are better $p$-values for building prediction sets in the iid model).
@@
We consider a Gaussian mixture model with $K=3$ components, where each component is bivariate normal. The centers for components 1,2,3 are, respectively, (0,0), (SNR,0), (SNR,SNR). So the overlap between components is larger as SNR decreases. We consider the case of balanced classes in the calibration sample, as well as the case of unbalanced classes where the null class is much larger than the others. Specifically, the balanced case has class probabilities 0.33, 0.33, and 0.34, and the unbalanced case has class probabilities 0.15, 0.10, and 0.75 (the null class). In the balanced case, we consider the iid setting where the test sample has the same class probabilities as the calibration sample (depicted in Figure 5 for an SNR value of 3), as well as class conditional setting where the test sample has class probabilities 0.2,0.2, and 0.6 (the null class), so the label shift is large. We estimate the probability of being in each class with a support vector classifier implemented by the e1071 R package Meyer et al., (2023).
@@
Figure 6 shows the FCR and resolution-adjusted power of all the procedures considered. As expected, the classic conformal procedure does not control the FCR for any data generating model (it uses the full-calibrated $p$-values in the iid setting, and the class-calibrated $p$-values in the class conditional setting). All other procedures control the FCR.
@@
For the iid model, InfoSCOP on full calibrated $p$-values has better power than the alternatives for prediction sets excluding a null class. Its advantage over InfoSP is primarily due to the fact that after pre-processing, almost all test examples are non-null, as illustrated in Figure 5 for a single data generation. The differences between the procedures are qualitatively the same, but even greater, when the overlap between components is larger, see Figure 13 in the SM. This is because when the overlap with the null class is large, after pre-processing only examples with better scores are considered, as illustrated in Figure 14 in the SM. For completeness, we also provide InfoSP on class calibrated $p$-values in the iid setting, to demonstrate numerically the potentially large power advantage from using the full calibrated $p$-values over the class calibrated $p$-values.
@@
For the class conditional model, InfoSP has lower power than classic conformal, but the power is reasonable. In Figure 9 in the SM we compare it to two procedures that weigh the classes according to their estimated relative frequencies.
@@
Figure 6: Selecting prediction sets excluding a null class in a classification setting. FCR (top row), and resolution-adjusted power (bottom row) versus SNR. The iid setting in columns 1 and 2, with balanced classes and unbalanced classes, respectively. The class conditional setting in column 3, with a large label shift: the class probabilities are equal in the calibration sample and 0.2,0.2, and 0.6 (the null class) in the test sample. The number of data generations was 2000, 1000 data points were used for training, and $n=m=500$. See details of the data generation in § 5.3.
@@ 5.4 Selecting non-trivial prediction sets
Suppose the analyst is interested in reporting prediction sets that are not equal to $[K]$ (see first item of Examples 1.2 and 2.2). In that case, we argue that InfoSP has an FCR close to $\alpha$ in the iid model. Intuitively, this comes from the selection rule $\mathcal{S}=\mathrm{BH}(\mathbf{\overline{p}})$ which is such that
${\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}}),i\in\mathcal{S}\}}={\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}})\}}.$
It means that it is not possible to fail to cover at the adjusted level without being selected. This is not the case in other selection rules, e.g., excluding a null class, where it is possible that the true class label is not covered at the adjusted level even if the example is not selected (thus implying that the adjusted level is conservative, since for FCR control we guard against non-coverage at the adjusted level for all examples). We formalize fully the argument for $K=2$ in the following result.
@@
In the iid classification model with $K=2$, consider the non-trivial informative subset collection =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq 1\}$ and assume that the score functions satisfy Assumption 1 with $\sum_{k\in[K]}S_{k}(x)=1$ and $S_{k}(x)\geq 0$. Let $p_{0}$ be the probability that $S_{Y_{i}}(X_{i})$ is the maximum score $\max\{S_{1}(X_{i}),S_{2}(X_{i})\}$. Then we have
$\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}),P_{X,Y})=\alpha(1-(1-p_{0})^{n+1}),$
provided that $(n+1)\alpha/m$ is an integer.
@@
The proof is given in § C.4, which also shows that $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$ coincides with the procedure of Zhao and Su, (2023) for $K=2$. In typical applications (where classes are not very well separated) the value of $(1-(1-p_{0})^{n+1})$ is close to one, which means that the FCR of our procedure should be close to $\alpha$, at least for $K=2$.
@@
To complement our theoretical result, we provide numerical results in Figure 7. Clearly, InfoSCOP is not improving InfoSP for non-trivial selection. InfoSP has the best power, with InfoSCOP a close second, on full calibrated $p$-values. InfoSP on class calibrated $p$-values has much lower power in the unbalanced setting. As expected, the classic conformal procedure does not control the FCR and the level for InfoPS is about $0.05$ for a range of SNR values. All other procedures control the FCR.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
6 Application to directional FDR control

### v3

@@ 5 Application to classification
We consider the classification case $\mathcal{Y}=[K]$, for both the iid model and the class-conditional model. Importantly, in classification, any selective prediction set procedure $\mathcal{R}=(\mathcal{C}_{n+i})_{i\in\mathcal{S}}$ is post-processed by setting, for $i\in\mathcal{S}$, $\mathcal{C}_{n+i}=\arg\min_{k\in[K]}\{S_{k}(X_{n+i})\}$ whenever $\mathcal{C}_{n+i}=\emptyset$ (that is, if empty take the smallest non-conformity score). Clearly, this operation can only decrease the FCP while it can only increase the adjusted power, so it should always be preferred in the classification case. In this paper, InfoSP and InfoSCOP always refer to the post-processed procedures in the classification case.
@@ 5.1 Choosing the appropriate $p$-value collection in classification
Thus, for the iid model, we can in principle use either full-calibrated $p$-values or class-calibrated $p$-values. In the applications we consider next, it appears that using full-calibrated $p$-values in InfoSP is best. We support this claim by theory for non-trivial classification in § 5.3, and by numerical experiments in § 5.2, § 5.3, and § J. We note that for the initial selection step in InfoSCOP, class-calibrated $p$-values can be useful, as demonstrated in § 5.2.
For the class-conditional model, there can be a label shift from the calibration to the test sample. So the full-calibrated $p$-values are not valid, and class-calibrated $p$-values must be used. In § D we consider more generally weighted class-calibrated $p$-values, where the weights are functions of estimators of the proportion of labels in each class in the test sample.
@@ 5.2 An illustrative example: prediction sets excluding a null class
Suppose the analyst is interested in reporting prediction sets that exclude a null class, say class $y_{0}=1$ (see first items of Examples 1.2 and 2.2). We consider the following novel procedures, in addition to the naive procedure using the classic conformal procedure, that reports $\mathcal{C}^{\alpha}_{n+i}$ only if $\mathcal{C}^{\alpha}_{n+i}$ does not intersect the null class: first, InfoSP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$. Second, InfoSCOP on full-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSCOP}}}(\mathbf{\overline{p}})$, with initial selection step $\mathcal{S}^{(0)}\subset[r+1,n+m]$ being the BH procedure applied to the class-calibrated adaptive $p$-value family $(\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}},i\in[r+1,n+m])$, given by $\tilde{p}^{(1)}_{i,{\tiny\mbox{adapt}}}=\hat{\pi}_{1}\tilde{p}^{(1)}_{i},\>i\in[r+1,n+m],$ using $\{(X_{j},Y_{j}),j\in[r]\}$ and $\{(X_{j},Y_{j}),j\in[r+1,n+m]\}$ as calibration and test samples, respectively, and with calibrated-based estimator $\hat{\pi}_{1}=(r+1)^{-1}\big{(}\sum_{i=1}^{r}{\mathbf{1}\{Y_{i}=1\}}+1\big{)}$. Third, InfoSP on class-calibrated $p$-values, denoted by $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}})$. Note that InfoSCOP above uses class-calibrated $p$-values for initial selection (because these are better to detect examples from the non-null class), and full-calibrated $p$-values on the selected examples from $[r+1,n+m]$ in the second step (because these are better $p$-values for building prediction sets in the iid model).
@@
We consider a Gaussian mixture model with $K=3$ components, where each component is bivariate normal. The centers for the three components are (0,0), (SNR,0), and (SNR,SNR). So the overlap between components is larger as SNR decreases. We consider the case of balanced classes in the calibration sample, as well as the case of unbalanced classes where the null class is much larger than the others. Specifically, the balanced case has class probabilities 0.33, 0.33, and 0.34, and the unbalanced case (depicted in Figure 11 for an SNR value of 3) has class probabilities 0.15, 0.10, and 0.75 (the null class). In the balanced case, we consider the iid setting where the test sample has the same class probabilities as the calibration sample, as well as class-conditional setting where the test sample has class probabilities 0.2,0.2, and 0.6 (the null class), so the label shift is large. We estimate the probability of being in each class with a support vector classifier implemented by the e1071 R package Meyer et al., (2023).
@@
Figure 4 shows the FCR and resolution-adjusted power of all procedures considered. As expected, the classic conformal procedure does not control the FCR for any data generating model (it uses the full-calibrated and class-calibrated $p$-values in the iid and class-conditional setting, respectively). All other procedures control the FCR.
@@
For the iid model (Figure 4 left and middle columns), InfoSCOP on full-calibrated $p$-values has better power than the alternatives for prediction sets excluding a null class. Its advantage over InfoSP is primarily due to the fact that after pre-processing, almost all test examples are non-null, as illustrated in Figure 11 in the SM for a single data generation. The differences between the procedures are qualitatively the same, but even greater, when the overlap between components is larger, see Figure 13 in the SM. This is because when the overlap with the null class is large, after pre-processing only examples with better scores are considered, as illustrated in Figure 14 in the SM. For completeness, we also provide InfoSP on class-calibrated $p$-values in the iid setting, to demonstrate numerically the potentially large power advantage from using the full-calibrated $p$-values over the class-calibrated $p$-values.
@@
For the class-conditional model (Figure 4 right column), InfoSP has lower power than classic conformal, but the power is reasonable. In Figure 6 in the SM we compare it to two procedures that weigh the classes according to their estimated relative frequencies.
@@
Figure 4: Selecting prediction sets excluding a null class in a classification setting. FCR (top row), and resolution-adjusted power (bottom row) versus SNR. The iid setting in columns 1 and 2, with balanced classes and unbalanced classes, respectively. The class-conditional setting in column 3, with a large label shift: the class probabilities are equal in the calibration sample and 0.2,0.2, and 0.6 (the null class) in the test sample. The number of data generations was 2000, 1000 data points were used for training, and $n=m=500$. See details of the data generation in § 5.2.
@@ 5.3 Selecting non-trivial prediction sets
Suppose the analyst is interested in reporting prediction sets that are not equal to $[K]$ (see first item of Examples 1.2 and 2.2). In that case, we argue that InfoSP has an FCR close to $\alpha$ in the iid model. Intuitively, this comes from the selection rule $\mathcal{S}=\mathrm{BH}(\mathbf{\overline{p}})$ which is such that ${\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}}),i\in\mathcal{S}\}}={\mathbf{1}\{Y_{n+i}\notin\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}(\mathbf{\overline{p}})\}}.$ It means that it is not possible to fail to cover at the adjusted level without being selected (because otherwise the prediction set is trivial). This is not the case for other selection rules, e.g., excluding a null class, where it is possible that the true class label is not covered at the adjusted level even if the example is not selected (thus implying that the adjusted level is conservative, since for FCR control we guard against non-coverage at the adjusted level for all examples). We formalize fully the argument for $K=2$ in the following result.
@@
In the iid classification model with $K=2$, consider the non-trivial informative subset collection =$\mathcal{I}=\{C\subset[K]\>:\>|C|\leq 1\}$ and assume that the score functions satisfy Assumption 1 with $\sum_{k\in[K]}S_{k}(x)=1$ and $S_{k}(x)\geq 0$. Let $p_{0}$ be the probability that $S_{Y_{i}}(X_{i})$ is the maximum score $\max\{S_{1}(X_{i}),S_{2}(X_{i})\}$. Then if $(n+1)\alpha/m$ is an integer, we have $\mathrm{FCR}(\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}}),P_{X,Y})=\alpha(1-(1-p_{0})^{n+1})$.
@@
The proof is given in § E.4, which also shows that $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\overline{p}})$ coincides with the procedure of Zhao and Su, (2023) for $K=2$. In typical applications (where classes are not very well separated) the value of $(1-(1-p_{0})^{n+1})$ is close to one, which means that the FCR of our procedure should be close to $\alpha$, at least for $K=2$.
@@
To complement our theoretical result, we provide numerical results in Figure 12. InfoSP has the best power, with InfoSCOP a close second, on full-calibrated $p$-values. InfoSP on class-calibrated $p$-values has much lower power in the unbalanced setting. As expected, the classic conformal procedure does not control the FCR and the level for InfoSP is about $0.05$ for a range of SNR values. All other procedures control the FCR.
6 Informative prediction sets for 3 classes of animals
In this section, we illustrate the performance of our methods on real data. We use the image dataset CIFAR-10 (https://www.cs.toronto.edu/~kriz/cifar.html), which consists of 60000 32x32 colour images in 10 classes, with 6000 images per class. We restrict the analysis to 3 classes: birds, cats and dogs. The code used for these experiments can be found at https://github.com/arianemarandon/infoconf.
We consider 4 scenarios: in the iid setting, non-trivial classification (scenario a) and non-null classification (scenario b); in the class-conditional setting, non-trivial classification (scenario c) and non-null classification (scenario d). The null class is the bird class in scenarios b and d. By default, the classes are in equal proportions. We introduce a label shift between the calibration sample and test sample for the class-conditional settings, by modifying the classes proportions: of the calibration sample to 20% for the bird class versus 40% for the cat/dog class in scenario c; of the test sample to 50% for the bird class versus 25% for the cat/dog class in scenario d. In each experiment, the test size is $m=1000$ and the calibration size is $n=5000$.
In the iid settings (scenarios a and b), we evaluate the procedures InfoSP and InfoSCOP with full-calibrated $p$-values and in the class-conditional settings (scenarios c) and d)) we evaluate InfoSP with class-calibrated $p$-values, each being also compared with classical conformal prediction (denoted by CC). For all procedures the non-conformity score is $S_{y}(x)=1-\pi_{y}(x)$ where $\pi_{y}(x)$ is an estimator of the probability that the class of $x$ is $y$ and is learned using a convolutional neural network (CNN) with 2 convolutional layers, one pooling layer, and 3 fully-connected layers, trained for 20 epochs with a learning rate of 0.01 on a sample size of 5000. To assess the power of the methods, in addition to the resolution-adjusted power, we plot the selection rate (SR) of the procedures, defined as the average proportion of informative prediction sets returned, and the average size of the informative prediction sets.
In each setting, the FDR and each power metric for the methods are evaluated by using 100 runs and the results are reported in Figure 5 scenarios a and b and in Figure 15 scenarios c and d for $\alpha=0.1$. The conclusions are qualitatively similar to the experiments of § 5: in all settings, classical conformal prediction yields an FCR that severely exceeds the marginal level with an inflation of about 50%. By contrast, our procedures InfoSP and InfoSCOP control the FCR at the target nominal level, both without and with label shift. In terms of power, concerning the iid settings, pre-processing is not useful for non-trivial classification as expected and the performances of InfoSP and InfoSCOP are similar with an FCR close to the nominal level for both in that case. For the non-null classification task, InfoSP is conservative while InfoSCOP is more powerful and displays an FCR close to $\alpha$. When there is label shift, in the case of non-trivial classification InfoSP displays an FCR close to $\alpha$. In the non-null classification case, however, the label shift increases the difficulty of the task in the sense that for a fixed selection, under-covering the null class in the test sample results in more false informative sets. Hence, the power is low in that case. Finally, in all scenarios, our procedures InfoSP and InfoSCOP output a lower number of informative sets compared to CC, but this is necessary in order to control the FCR. The average size of the prediction sets that are informative is comparable.
a) Non-trivial classification
b) Non-null classification
Figure 5: FCR, average size of the selected, SR, and resolution-adjusted power for the methods, for $\alpha=0.1$.
7 Conclusion and discussion
Our methods are very general, and they are relevant to applications in classification and regression. We showed examples in § 4-§ 6 for informative subsets of interest. We used common scores for the examples, but many other scores from the literature can be used with our suggested procedures InfoSP and InfoSCOP as long as Assumption 1 is satisfied. In addition, transfer learning scores can also be handled by our theory (see Assumption 6 in SM). This is known to greatly improve the conformal inference when there is a domain shift between the learning sample and the calibration+test samples (Courty et al.,, 2017; Gazin et al.,, 2024) and developing specific application cases for the latter is of interest for future investigations.
For the iid model, InfoSCOP improves over InfoSP in all considered examples, except when selecting non-trivial prediction subsets in classification (Example 1.2 item 2), for which we establish that InfoSP almost exhausts the FCR level (see Proposition 5.1for $K=2$). The procedure InfoSCOP splits the calibration sample in order to apply an efficient initial selection step on part of the calibration sample and the test sample. There are many ways to perform the initial selection. The choice is important because it defines the pre-processed $p$-values (22) that can be seen as $p$-values “conditionally on being selected”. Different ways of choosing $S^{(0)}$ have been investigated: trying to rule out all the examples in null class (§ 5.2) or trying to mimic the $\mathrm{BH}(\mathbf{q})$ selection that will be applied at the second stage to reduce the selection effect (§ 4). Finding an optimal way of calibrating $S^{(0)}$ is an interesting avenue for future research.
We provided a class-conditional variant of InfoSP, with class-conditional guarantees. We proved that our strategy can be followed in the case where the classes of the calibration and test samples are arbitrary fixed, even when the class proportions in the calibration are very different than in the test. The main point is that working with the class-calibrated $p$-value collection allows to maintain the FCR control in this strong sense. In § D we suggest additionally weighted procedures, that incorporate the estimated class proportions. These procedures are not necessarily more powerful than InfoSP, and further research is needed in order to make recommendations about when to use the weighted procedures, and about weight adaptation to the specifics of the data.
Acknowledgments
The authors would like to thank two anonymous referees and an associate editor for helpful comments. The authors acknowledge grants ANR-21-CE23-0035 (ASCAI) and ANR-23-CE40-0018-01 (BACKUP) of the French National Research Agency ANR, the Emergence project MARS of Sorbonne Université, and Israeli Science Foundation grant no. 2180/20.
Appendix A Connections to existing works
For the class-conditional model, we can view $(Y_{n+i})_{i\in[m]}$ as fixed. Thus, informative prediction sets can be viewed as informative confidence sets for parameters. This has been considered in a particular setting by Weinstein and Yekutieli, (2020). They considered building confidence intervals only for the selected parameters that will be sign determining. They showed that if the test statistics are independent and the confidence intervals satisfy some monotonicity properties, then the FCR can be controlled. Their theoretical framework is different than the one we consider, but their approach of selecting only sign-determining confidence intervals is very similar to ours, of selecting only informative prediction sets when informativeness is defined by sign determination. Moreover, this approach has been considered in Weinstein and Ramdas, (2020) with the broader scope of only reporting confidence intervals if they are “localizing” appropriately the true parameter in the sense that the confidence interval is entirely contained in one element of a pre-specified partition of the $\mathcal{Y}$ space. They investigate this task in the online setting where the sequence of unknown parameters is fixed, and at each time step an independent observation is observed for the corresponding parameter, which is substantially different than our batch setting where the conformal $p$-values are dependent and the outcome may be random. We show in Remark 2.3 that our informativeness theory covers their localizing notion in our setting. Next, we discuss inspiring works connected to ours that assume the iid model.
Zhao and Su, (2023) suggested procedures for average error control in multi-class classification, so $\mathcal{C}_{n+i}$ are singletons. Since their procedure only reports a single class for each selected example, ambiguous examples will not be selected. However, in most classification tasks, there are examples whose true class is difficult to determine, yet it is possible to narrow down the possible set of classes (Sadinle et al.,, 2019). We suggest procedures that produce $\mathcal{C}_{n+i}$ that are not necessarily singletons for $K>2$. For $K=2$, their procedure coincides with an instance in our framework, see details in § 5.3. However, for $K>2$, while our suggestion as well as their suggestion provides level $\alpha$ FCR control, we select more examples, and although the prediction sets may be at a coarser resolution than singletons, they are still informative since they narrow down the possible set of classes. We note that for $K>2$, we can recover their procedure if we define as informative only prediction sets of size one, since then their procedure coincides with InfoSP for the iid model.
Jin and Candes, (2023) addressed the problem of discovering outcomes with values above a threshold. So $\mathcal{C}_{n+i}$ is of the form $(c_{i},\infty)$ for predefined $(c_{i})_{i\in[m]}$. They cast the problem as that of testing the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$. In § I.2, it is demonstrated that by defining $Y_{n+i}\leq c_{i}$ as uninformative, we can complement the discoveries of Jin and Candes, (2023) with one-sided prediction intervals, while providing the same false discovery rate guarantee on the selected. Moreover, we show how to obtain two-sided prediction intervals for the informative examples in § 4.1.
Bao et al., (2024) considered the regression framework. Their first result (Proposition 1 therein) is to prove that for selection rules that do not depend on the calibration sample, classic conformal prediction intervals at level $\alpha|\mathcal{S}|/m$ for the $\mathcal{S}$ selected examples (i.e., the correction factor $|\mathcal{S}|/m$ for selection suggested in Benjamini and Yekutieli,, 2005), provide level $\alpha$ FCR control. For our purpose of informative selection, this result is not useful because informative selection involves all conformal $p$-values and therefore involves the calibration sample in a specific way. We need a different set of assumptions that are detailed in our novel Theorem C.1.
Bao et al., (2024) further argue that the resulting prediction intervals, $\big{(}\mathcal{C}^{\alpha|\mathcal{S}|/m}_{n+i}\big{)}_{i\in\mathcal{S}}$ are too wide. They suggest a novel approach that performs selection on both the calibration set and test set, and then constructs $\alpha$ level conformal prediction intervals for the selected test candidates using the conditional empirical distribution obtained by the post-calibration set. For exchangeable selection rules, they show that the FCR is controlled at level $\alpha$. Their selection process cannot guarantee that all the constructed prediction intervals are of interest to the analyst. For example, for predicting the affinity of drug-target pairs, the analyst may not be interested in pairs with affinity below, say, $y_{0}$ (the case in item 1 of Example 1.1). Using the novel approach of Bao et al., (2024), prediction intervals will be constructed following the selection of calibration and test examples for which the predicted affinity from the machine learning algorithm is above a selection threshold. They require that the selection procedure be a thresholding procedure of the scores $S_{Y_{j}}(X_{j})$ with a threshold $\tau$ that is either independent of the calibration sample (their Proposition 1), or exchangeable with respect to both calibration and test samples (their Theorem 1). Some of these prediction intervals may include $y_{0}$, and thus be useless for the analyst (an illustration is given in § I.4 in SM, see Figure 10 therein). However, it is not possible to additionally select only the examples with prediction intervals above $y_{0}$, since after performing this additional selection, the prediction intervals of Bao et al., (2024) on the selected no longer have an $\alpha$ level FCR guarantee. Our procedures for regression thus complement the work of Bao et al., (2024) when the focus is that all the prediction intervals eventually constructed are informative.
In a very recent work [^4] This work appeared when we were in the final stage of writing, our work has been done independently., Jin and Ren, (2024) generalize the work of Bao et al., (2024), by considering more general selection rules for finite-sample exact coverage conditional on the unit being selected. Their conditional guarantee is achieved by a careful swapping argument which identifies for each selection rule, the appropriate subset of the calibration examples for each example from the test sample. Their conditional error guarantee implies FCR control under some conditions. The control is valid for any selection rule that is exchangeable within the calibration sample, which is a weaker assumption than concordance. Hence, their selection may be based on some notion of informativeness. However, since the prediction sets should be inflated after applying the selection, the latter are not necessary informative according to our definition. For example, after selecting by a multiple testing procedure on the family of null hypotheses $\{Y_{n+i}\leq c_{i},i\in[m]\}$ in the setting of Jin and Candes, (2023), their prediction sets may include the $c_{i}$’s for some of the discoveries. We suggest procedures where selection and construction of prediction sets are inseparable, since we require that each selected prediction set be informative (along with the requirement that FCR $\leq\alpha$).
Appendix B Application to directional FDR control